an object projected vertically up from top of a tower took 5 seconds to reach the ground. if the average velocity of the object is 5 metre per second, find its average speed. given gravity is equal to 10 metre per second.

Grade:9

1 Answers

Eshan
5 years ago
Let the height of tower behand velocity with which ball is thrown up bev.

Equation of motion for the ball’s total flight-

-h=ut+gt^2/2 \implies -h=5u-5t^2

Also the average velocity is given by\dfrac{h}{t}=5m/s\implies h=25m
Solving the two equations gives h=25m, u=20m/s

Maximum height attained by ball from the top of tower=\dfrac{u^2}{2g}=20m
Hence the average speed attained=\dfrac{20m+20m+25m}{5}m/s=13 m/s